OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5

OpenStudy (anonymous):

2+3 -2 +me +you =love

OpenStudy (anonymous):

Find f(4) and g(4) and subtract them.

OpenStudy (anonymous):

f(4)-g(4) f(x) = 2x − 2 and g(x) = 3x − 1 2(4)-2 - 3(4)-1 8-2 - 12-1 6-11 -5